OpenStudy (anonymous):

if 3 is one side and 5 is another side of a triangle, what could the third side be? dont i just add these two sides together? meaning the third side would 8?

OpenStudy (anonymous):

please guys this is an easy question.. just wanna make sure

OpenStudy (anonymous):

Is this a right angle triangle?

OpenStudy (anonymous):

all the question says is "nancy cuts a piece of cardboard in the shape of a triangle"

OpenStudy (anonymous):

@purplec16 -all the question says is "nancy cuts a piece of cardboard in the shape of a triangle"

OpenStudy (anonymous):

Well, am pretty sure you don't add them give me a minute to refresh my memory I haven't done this in years..

OpenStudy (anonymous):

i am pretty dump with this so i have no idea

OpenStudy (anonymous):

dumb** lol

OpenStudy (anonymous):

No you're not... and I really don't remember @dpalnc

OpenStudy (anonymous):

i didnt think this question was gonna be this hard hehe

OpenStudy (anonymous):

It's not hard I just... don't remember but I'm pretty sure you don't add them...

OpenStudy (ganpat):

Generally area of triangle is given by, A = 1/2 x base x height or Area = 1/2 x one side x other side x sin of angle opposite opposite to unknown side.. Look if it can help !!!

OpenStudy (lgbasallote):

@aria11 what you need to know about triangle is that the sum of two sides should be greater than the third side so let's say the third side is x 5 + 3 > x so 8 > x so the third side could be 8 and below but wait remember i said sum of two sides must be greater than the third so the sum of x and 3 should be greater than 5 x + 3 > 5 x > 2 so the third side should be greater than 2 but less than 8 so 2<x<8 got that?

OpenStudy (anonymous):

hi use pythagourus theorm.

OpenStudy (anonymous):

She doesn't know if it's a right angle triangle so she can't use pythagorean @muhammad9t5

OpenStudy (anonymous):

find angle between them first.

OpenStudy (anonymous):

or use law of cosine.

OpenStudy (anonymous):

friend i thinnk you also should have an angle then you can solve it easily.

OpenStudy (ganpat):

There is more sine called sine rule, you may use that if you know all angles as well.. a/SinA = b/ SinB = c/ SinC

OpenStudy (anonymous):

hi now i am on your problem. Solution is let us consider that 3 is adjacent and 5 is hypotenuse therefore cosx=adjacent/hyp cosx=3/5 cosx=0.6 x=\[\cos^{-1} 0.6 \] x=53.13 deg wae have found angle now we have to find the side \[c ^{2}=a ^{2}+b ^{2}-2abcocx\] \[c ^{2}=3^{2}+5^{2}-2\times3\times5\times0.6\] \[c ^{2}=9+25-18\]\[c ^{2}=34-18\] \[c ^{2}=16\] c=4 i hope it works.